# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 5.4 Landau Damping and Transit-Time Damping
For simplicity in this analysis, we will assume the isotropic temperature case for a proton-electron plasma with linear wave properties given by the isotropic linear Vlasov-Maxwell dispersion relation (22); extending these calculations to allow for anisotropic temperatures can be done simply following the same logical procedure and using the linear Vlasov-Maxwell dispersion relation for anisotropic temperatures given by (21). Furthermore, justified by observations in solar wind turbulence that incompessible Alfvénic fluctuations energetically dominate the turbulence (Belcher & Davis, 1971; Tu & Marsch, 1995; Alexandrova et al., 2008; Bruno & Carbone, 2013), we will focus on the linear collisionless damping rates of the Alfvénic fluctautions, including both Alfvén waves in the MHD limit at $k_{\perp}\rho_{i}\ll 1$, and kinetic Alfvén waves at $k_{\perp}\rho_{i}\gtrsim 1$. Consideration of the damping of compressible turbulent fluctuations, specifically fast and slow magnetosonic fluctuations at $k_{\perp}\rho_{i}\ll 1$ (Howes et al., 2012a), is easily performed using the same procedure with the linear dispersion relations for these alternative wave modes [^7] Although, it is important to keep in mind that the recent identified nonlinear effect of anti-phase-mixing in turbulent plasmas can suppress the collisionless damping predicted by linear kinetic theory for the compressible wave modes in the inertial range of scales at $k_{\perp}\rho_{i}\ll 1$ (Schekochihin et al., 2016; Parker et al., 2016)..
@@
where $\overline{\omega}\equiv\omega/k_{\parallel}v_{A}$. For Alfvén waves, this normalized frequency $\overline{\omega}$ is well approximated by (Howes et al., 2014)
$\overline{\omega}(k_{\perp}\rho_{i},\beta_{i},\tau)\equiv\frac{\omega}{k_{\parallel}v_{A}}=\sqrt{1+\frac{(k_{\perp}\rho_{i})^{2}}{\beta_{i}+2/(1+1/\tau)}},$ (25)
@@
Whether damping by the Landau-resonant mechanisms is weak or strong primarily depends on where the parallel wave phase velocity falls within the equilibrium particle velocity distribution, as quantified by the expressions in (24). The application of the weak-growth-rate approximation to the kinetic theory for Landau damping (Krall & Trivelpiece, 1973) indicates that the linear damping rate is proportional to the slope of the distribution at the resonant velocity, as depicted in Fig. 6. For phase velocities that fall deep in the core of the velocity distribution $\omega/k_{\parallel}\ll v_{ts}$ (blue dotted) where the slope of $f_{s}(v_{\parallel})$ is flat, the damping is weak because nearly equal numbers of particles with $v_{\parallel}<\omega/k_{\parallel}$ are accelerated by the wave as particles with $v_{\parallel}>\omega/k_{\parallel}$ are decelerated by the wave, leading to little net energy exchange. Similarly, for phase velocities far out in the tail of the velocity distribution $\omega/k_{\parallel}\gg v_{ts}$ (black dashed), there are few particles to resonate with the waves and the slope of $f_{s}(v_{\parallel})$ is also flat, so the damping is weak. Only if the parallel phase velocity falls in the steep region of the velocity distribution at $\omega/k_{\parallel}\sim v_{ts}$ (red solid) is there a significant net energy transfer to the particles: more particles with $v_{\parallel}<\omega/k_{\parallel}$ are accelerated by the wave than those particles with $v_{\parallel}>\omega/k_{\parallel}$ are decelerated by the wave, so that particles gain net energy at the expense of the wave.
@@
The ultimate goal is to characterize the turbulent plasma heating— specifically $Q_{i}/Q_{e}$, $Q_{\perp,i}/Q_{i}$,and $Q_{\perp,e}/Q_{e}$—due to Landau damping and transit-time damping as a function of the fundamental dimensionless parameters of turbulence. To accomplish this goal, it is necessary to integrate the damping rates over the full range of the turbulent cascade, as discussed in §5.2 and illustrated in Figures 4 and 5. This integration over all scales effectively eliminates the dependence of the damping mechanisms on the components of the wavevector $k_{\perp}\rho_{i}$ and $k_{\parallel}\rho_{i}$, leaving only the desired dependence on the plasma and turbulence parameters enumerated in Tables 2 and  3.
@@
With the scale dependence on $k_{\perp}\rho_{i}$ eliminated, the ion Landau damping (iLD) and ion transit-time damping (iTTD) rates depend primarily on $\beta_{i}$, and the electron Landau damping (eLD) and electron transit-time damping (eTTD) rates depend on $\beta_{i}$, $\tau$, and $\mu$, as shown by (24). Note that the dependence of eLD and eTTD on $\tau$ and $\mu$ primarily arises from those parameters controlling the scaling of the ion to electron Larmor radius scales, $\rho_{i}/\rho_{e}\sim(\tau\mu)^{1/2}$.
@@
It is worthwhile pointing out that the physics of collisionless damping via the Landau resonance is fully captured in the gyrokinetic approximation, with the linear wave frequencies and damping rates both linearly proportional to $k_{\parallel}$, as shown in (23). Thus, the ratio of the damping-to-cascade rate [^8] Here we have adopted the condition of criticial balance for strong plasma turbulence $\omega_{nl}\simeq\omega$ (Goldreich & Sridhar, 1995; Mallet et al., 2015)., $\gamma/\omega_{nl}\simeq\gamma/\omega$, is formally independent of $k_{\parallel}$, leaving the expected parameter dependence on the plasma parameters $\beta_{i}$ and $\tau$, along with the mass ratio $\mu$.
@@
Note that dependence of the linear gyrokinetic dispersion relation on $(k_{\perp}\rho_{i},\beta_{i},\tau)$ in (23) is consistent with the results of our analysis of the parameter dependence of the Landau-resonant damping mechanisms. The linear gyrokinetic dispersion relation also determines the amplitudes and phases of the key electromagnetic fields $E_{\parallel}$ and $\delta B_{\parallel}$ that mediate the energization by Landau damping and transit-time damping. As done in the H10 turbulent heating model, the linear dispersion relation can simply be used to calculate directly the linear collisionless damping rates as a function of $(k_{\perp}\rho_{i},\beta_{i},\tau)$, but these linear damping rate predictions may be modified under strongly turbulent plasma conditions. However, field-particle correlation analyses of numerical simulations (Klein et al., 2017, 2020; Huang et al., 2024) and spacecraft observations (Chen et al., 2019; Afshari et al., 2021) of plasma turbulence have clearly shown that Landau damping and transit-time damping do indeed play a role in the damping of strong plasma turbulence.
@@
One final step is to estimate how the ion-to-electron heating ratio $Q_{i}/Q_{e}$ depends on the remaining key parameters $\beta_{i}$, $\tau$, and $\mu$. As illustrated in Fig. 5(c) and Fig. 7, the ion damping rate $\gamma_{i}$ peaks at scales $k_{\perp}\rho_{i}\sim 1$, and this Landau-resonant damping onto the ions is essentially independent of the ratio ion-to-electron temperature ratio. The electron damping rate $\gamma_{e}$, on the other hand, increases monotonically as the perpendicular wavevector increases to the electron Larmor radius scale $k_{\perp}\rho_{e}\rightarrow 1$, eventually yielding a normalized collisionless damping rate that is sufficiently strong to terminate the turbulent cascade with $\gamma_{e}/\omega\rightarrow 1$, as shown in Fig. 7. Thus, any energy in the turbulent cascade that passes beyond the ion kinetic scales at $k_{\perp}\rho_{i}\sim 1$ will ultimately be transferred to the electrons. Since the ion damping rate depends dominantly on $\beta_{i}$ with very weak dependence on $\tau$, yielding $Q_{i}(\beta_{i})$ to lowest order, this means that the ratio $Q_{i}/Q_{e}$ is effectively determined by the ion plasma beta alone given a total turbulent energization rate that eventually must balance the total cascade rate $Q\sim\varepsilon_{0}$, since it can be shown that $Q_{e}/Q_{i}=\varepsilon_{0}/Q_{i}(\beta_{i})-1$. This dominant dependence of the Landau-resonant collisionless damping mechanisms on $\beta_{i}$ is consistent with existing turbulent heating models (Howes, 2010; Kawazura et al., 2019).
One final unexpected physical behavior of Landau damping and transit-time damping is also illustrated in Fig. 7. For isotropic Maxwellian velocity distributions of the ions and electrons, the imaginary component of the complex frequency, $\omega_{c}\equiv\omega-i\gamma$, determined by the linear dispersion relation is always negative (solid black curve), corresponding to a damping of the wave with $\gamma>0$. However, for this particular plasma parameter choice of $\beta_{i}=3$, $T_{i}/T_{e}=1$, and $m_{i}/m_{e}=1836$, the individual energy transfer rates to the ions through Landau damping mediated by $E_{\parallel}$ are actually negative over the range $k_{\perp}\rho_{i}\lesssim 2$ (red dotted), indicating a net transfer of energy from the ions to the wave fields when averaged over the full period (or, alternatively, over $2\pi$ in phase) of the Alfvén wave. But the contribution from transit-time damping mediated by $E_{\perp}$ (Huang et al., 2024; Howes et al., 2024) is positive and much larger (green long dashed), leading to a net damping of the wave ($\gamma_{i}>0$, thin blue) by the sum of the Landau-resonant interactions of ions with the Alfvén wave. Whether the wave-period-averaged energy transfer to ions is positive or negative depends on the relative phases of $E_{\parallel}$ and $E_{\perp}$ to the self-consistently determined ion current density $\mathbf{j}_{i}$. For the case of $\beta_{i}=3$ pictured in Fig. 7, at scales $k_{\perp}\rho_{i}\lesssim 2$ (red dotted), this relative phase $\delta\phi$ between $E_{\parallel}$ and $j_{\parallel,i}$ must fall in the range $\pi/2\leq\delta\phi\leq 3\pi/2$ (corresponding to wave growth at the expense of ion energy), while the relative phase between $E_{\perp}$ and $j_{\perp,i}$ must fall in the range $-\pi/2\leq\delta\phi\leq\pi/2$ (corresponding to wave damping and ion energy gain). Because each individual mechanism can possibly lead to negative energy transfer while the sum always leads to damping, it suggests that perhaps a separation of the particle energization rates by Landau damping and transit-time damping is not particularly important in unraveling the turbulent heating of the plasma (especially since both of these mechansims ultimately lead to energization of the parallel degree of freedom of the particles), so that only the summed effect of the Landau-resonant collisionless damping rates is important.
@@
Figure 7: Plot of the normalized damping rate $\gamma/\omega$ (black dashed) for the Alfvén and kinetic Alfvén wave over the perpendicular wavenumber range $10^{-2}\leq k_{\perp}\rho_{i}\leq 10^{2}$ for a plasma with $\beta_{i}=3$, $T_{i}/T_{e}=1$, and $m_{i}/m_{e}=1836$, where the complex eigenfrequency is given by $\omega_{c}=\omega-i\gamma$, so that $\gamma>0$ corresponds to wave damping and $\gamma<0$ to wave growth. The contributions from the ion damping rate $\gamma_{i}$ (thin blue) and electron damping rate $\gamma_{e}$ (thin magenta) are separately plotted, with this decomposition failing at strong damping rates with $\gamma/\omega\gtrsim 0.5$, occurring for $k_{\perp}\rho_{i}\gtrsim 40$. Ion damping due to Landau damping $\gamma_{i,LD}>0$ (red dashed) and transit-time damping $\gamma_{i,TTD}>0$ (green long dashed) are separated, showing a region of net energy transfer from ions to the wave for ion Landau damping with $\gamma_{i,LD}>0$ (red dotted) at $k_{\perp}\rho_{i}\leq 2$.

### v2

@@ 5.4 Landau Damping and Transit-Time Damping
For simplicity in this analysis, we will assume the isotropic temperature case for a proton-electron plasma with linear wave properties given by the isotropic linear Vlasov-Maxwell dispersion relation (22); extending these calculations to allow for anisotropic temperatures can be done simply following the same logical procedure and using the linear Vlasov-Maxwell dispersion relation for anisotropic temperatures given by (21). Furthermore, justified by observations in solar wind turbulence that incompessible Alfvénic fluctuations energetically dominate the turbulence (Belcher & Davis, 1971; Tu & Marsch, 1995; Alexandrova et al., 2008; Bruno & Carbone, 2013), we will focus on the linear collisionless damping rates of the Alfvénic fluctautions, including both Alfvén waves in the MHD limit at $k_{\perp}\rho_{i}\ll 1$, and kinetic Alfvén waves at $k_{\perp}\rho_{i}\gtrsim 1$. Consideration of the damping of compressible turbulent fluctuations, specifically fast and slow magnetosonic fluctuations at $k_{\perp}\rho_{i}\ll 1$ (Howes et al., 2012a), is easily performed using the same procedure with the linear dispersion relations for these alternative wave modes [^9] Although, it is important to keep in mind that the recently identified nonlinear effect of anti-phase-mixing in turbulent plasmas can suppress the collisionless damping predicted by linear kinetic theory for the compressible wave modes in the inertial range of scales at $k_{\perp}\rho_{i}\ll 1$ (Schekochihin et al., 2016; Parker et al., 2016; Meyrand et al., 2019)..
@@
where we define the dimensionless wave frequency normalized the the Alfvén wave frequency in the MHD limit,
$\overline{\omega}\equiv\frac{\omega}{k_{\parallel}v_{A}}$ (25)
For Alfvén waves, this normalized frequency $\overline{\omega}$ is well approximated by (Howes et al., 2014)
$\overline{\omega}(k_{\perp}\rho_{i},\beta_{i},\tau)=\sqrt{1+\frac{(k_{\perp}\rho_{i})^{2}}{\beta_{i}+2/(1+1/\tau)}},$ (26)
@@
Whether damping by the Landau-resonant mechanisms is weak or strong primarily depends on where the parallel wave phase velocity falls within the equilibrium particle velocity distribution, as quantified by the expressions in (24). The application of the weak-growth-rate approximation to the kinetic theory for Landau damping (Krall & Trivelpiece, 1973) indicates that the linear damping rate is proportional to the slope of the distribution at the resonant velocity, as depicted in Fig. 6. For phase velocities that fall deep in the core of the velocity distribution $\omega/k_{\parallel}\ll v_{ts}$ (blue dotted) where the slope of $f_{s}(v_{\parallel})$ is relatively flat, the damping is weak because nearly equal numbers of particles with $v_{\parallel}<\omega/k_{\parallel}$ are accelerated by the wave as particles with $v_{\parallel}>\omega/k_{\parallel}$ are decelerated by the wave, leading to little net energy exchange. Similarly, for phase velocities far out in the tail of the velocity distribution $\omega/k_{\parallel}\gg v_{ts}$ (black dashed), there are few particles to resonate with the waves and the slope of $f_{s}(v_{\parallel})$ is also quite flat, so the damping is weak. Only if the parallel phase velocity falls in the steep region of the velocity distribution at $\omega/k_{\parallel}\sim v_{ts}$ (red solid) is there a significant net energy transfer to the particles: more particles with $v_{\parallel}<\omega/k_{\parallel}$ are accelerated by the wave than those particles with $v_{\parallel}>\omega/k_{\parallel}$ are decelerated by the wave, so that particles gain net energy at the expense of the wave. Note that this qualitative picture of collisionless damping by the Landau resonance being controlled by where the parallel wave phase velocity falls within the distribution function is independent of the form of the equilibrium velocity distribution; thus, one can apply the same reasoning to estimate the strength of the Landau-resonant damping for other non-Maxwellian forms of the equilibrium velocity distribution, such as the kappa distributions often applied to model in situ measurements in space plasmas (Vasyliunas, 1968; Abraham-Shrauner & Feldman, 1977; Gosling et al., 1981; Lui & Krimigis, 1981; Armstrong et al., 1983; Summers & Thorne, 1991; Thorne & Summers, 1991; Summers et al., 1994; Livadiotis & McComas, 2013; Livadiotis et al., 2018).
@@
The ultimate goal is to characterize the turbulent plasma heating— specifically $Q_{i}/Q_{e}$, $Q_{\perp,i}/Q_{i}$,and $Q_{\perp,e}/Q_{e}$—due to Landau damping and transit-time damping as a function of the fundamental dimensionless parameters of turbulence. To accomplish this goal, it is necessary to integrate the damping rates over the full range of the turbulent cascade, as discussed in §5.2 and illustrated in Figures 4 and 5. This integration over all scales effectively eliminates the dependence of the damping mechanisms on the components of the wavevector $k_{\perp}\rho_{i}$ and $k_{\parallel}\rho_{i}$, leaving only the desired dependence on the plasma and turbulence parameters enumerated in Tables 3 and 4.
@@
With the scale dependence on $k_{\perp}\rho_{i}$ eliminated, the ion Landau damping (iLD) and ion transit-time damping (iTTD) rates depend primarily on $\beta_{i}$, and the electron Landau damping (eLD) and electron transit-time damping (eTTD) rates depend on $\beta_{i}$, $\tau$, and $\mu$, as shown by (24). Note that the dependence of eLD and eTTD on $\tau$ and $\mu$ primarily arises from those two parameters controlling the scale separation of the ion and electron Larmor radius scales, $\rho_{i}/\rho_{e}=(\tau\mu)^{1/2}$.
@@
It is worthwhile pointing out that the physics of collisionless damping via the Landau resonance is fully captured in the gyrokinetic approximation, with the linear wave frequencies and damping rates both linearly proportional to $k_{\parallel}$, as shown in (23). Thus, the ratio of the damping-to-cascade rate [^10] Here we have adopted the condition of critical balance for strong plasma turbulence $\omega_{nl}\simeq\omega$ (Goldreich & Sridhar, 1995; Mallet et al., 2015)., $\gamma/\omega_{nl}\simeq\gamma/\omega$, is formally independent of $k_{\parallel}$, leaving the expected parameter dependence on the plasma parameters $\beta_{i}$ and $\tau$, along with the mass ratio $\mu$, once the cascade is integrated over the perpendicular scales of the cascade.
@@
In addition to the fact that the dependence of the linear gyrokinetic dispersion relation on $(k_{\perp}\rho_{i},\beta_{i},\tau)$ in (23) is consistent with the results of our analysis of the parameter dependence of the Landau-resonant damping mechanisms, the linear gyrokinetic dispersion relation also determines the amplitudes and phases of the key electromagnetic fields $E_{\parallel}$ and $\delta B_{\parallel}$ that mediate the particle energization by Landau damping and transit-time damping. As done in the H10 turbulent heating model (see §4), the linear dispersion relation can simply be used to calculate directly the linear collisionless damping rates as a function of $(k_{\perp}\rho_{i},\beta_{i},\tau)$. It is possible, of course, that these linear damping rates may be modified under strongly turbulent plasma conditions, but modest quantitative changes to the effective damping rates are typically absorbed into the adjustable constant typically used in cascade models to specify the balance between the nonlinear cascade rate and collisionless damping rate (Howes et al., 2008a, 2011). However, field-particle correlation analyses of numerical simulations (Klein et al., 2017, 2020; Horvath et al., 2020; Conley et al., 2023; Huang et al., 2024) and spacecraft observations (Chen et al., 2019; Afshari et al., 2021) of plasma turbulence have clearly shown that Landau damping and transit-time damping do indeed play a role in the damping of strong plasma turbulence.
@@
One final step is to estimate how the ion-to-electron heating ratio $Q_{i}/Q_{e}$ depends on the remaining key parameters $\beta_{i}$, $\tau$, and $\mu$. As illustrated in Fig. 5(c) and Fig. 7, the ion damping rate $\gamma_{i}$ peaks at scales $k_{\perp}\rho_{i}\sim 1$, and numerical solutions of the Vlasov-Maxwell dispersion relation (not presented here) show that this Landau-resonant damping onto the ions is essentially independent of the ion-to-electron temperature ratio. The electron damping rate $\gamma_{e}$, on the other hand, increases monotonically as the perpendicular wavevector increases to the electron Larmor radius scale $k_{\perp}\rho_{e}\rightarrow 1$, eventually yielding a normalized collisionless damping rate that is sufficiently strong to terminate the turbulent cascade with $\gamma_{e}/\omega\rightarrow 1$, as shown in Fig. 7. Thus, any energy in the turbulent cascade that passes beyond the ion kinetic scales at $k_{\perp}\rho_{i}\sim 1$ will ultimately be transferred to the electrons. Since the ion damping rate depends dominantly on $\beta_{i}$ with very weak dependence on $\tau$, yielding $Q_{i}(\beta_{i})$ to lowest order, this means that the ratio $Q_{i}/Q_{e}$ is effectively determined by the ion plasma beta alone. This simplification arises because the total turbulent damping rate must eventually balance the total cascade rate, $Q=Q_{i}+Q_{e}\sim\varepsilon_{0}$, so it can be shown that $Q_{e}/Q_{i}=\varepsilon_{0}/Q_{i}(\beta_{i})-1$. This dominant dependence of the Landau-resonant collisionless damping mechanisms on $\beta_{i}$ is consistent with existing turbulent heating models for Alfvénic turbulence (Howes, 2010; Kawazura et al., 2019).
@@
Figure 7: Plot of the normalized damping rate $\gamma/\omega$ (black dashed) for the Alfvén and kinetic Alfvén wave over the perpendicular wavenumber range $10^{-2}\leq k_{\perp}\rho_{i}\leq 10^{2}$ for a plasma with $\beta_{i}=3$, $T_{i}/T_{e}=1$, and $m_{i}/m_{e}=1836$, where the complex eigenfrequency is given by $\omega_{c}=\omega-i\gamma$, so that $\gamma>0$ corresponds to wave damping and $\gamma<0$ to wave growth. The contributions from the ion damping rate $\gamma_{i}$ (thin blue) and electron damping rate $\gamma_{e}$ (thin magenta) are separately plotted, with the decomposition by species failing at strong damping rates with $\gamma/\omega\gtrsim 0.5$, occurring for $k_{\perp}\rho_{i}\gtrsim 40$. Ion damping due to Landau damping $\gamma_{i,LD}>0$ (red dashed) and transit-time damping $\gamma_{i,TTD}>0$ (green long dashed) are separated, showing a region of net energy transfer from ions to the wave by the Landau resonance with the ions yielding $\gamma_{i,LD}<0$ (red dotted) at $k_{\perp}\rho_{i}\leq 2$.
One final unexpected physical behavior of Landau damping and transit-time damping is also illustrated in Fig. 7. For isotropic Maxwellian velocity distributions of the ions and electrons, the imaginary component of the complex frequency, $\omega_{c}\equiv\omega-i\gamma$, determined by the linear dispersion relation is always negative (solid black curve), corresponding to a damping of the wave with damping rate $\gamma>0$. However, for the particular plasma parameter choices in Fig. 7 of $\beta_{i}=3$, $T_{i}/T_{e}=1$, and $m_{i}/m_{e}=1836$, the individual energy transfer rates to the ions through Landau damping [^11] The separation of the contributions of Landau damping and transit-time to the total collisionless damping rate for the Vlasov-Maxwell dispersion relation is detailed in Huang et al. (2024). mediated by $E_{\parallel}$ are actually negative over the range $k_{\perp}\rho_{i}\lesssim 2$ (red dotted), indicating a net transfer of energy from the ions to the wave fields when averaged over the full period (or, alternatively, over $2\pi$ in phase) of the Alfvén wave. But the contribution from transit-time damping mediated by $E_{\perp}$ (Huang et al., 2024; Howes et al., 2024) is positive and much larger (green long dashed), leading to a net damping of the wave ($\gamma_{i}>0$, thin blue) by the sum of the Landau-resonant ($n=0$) interactions of ions with the Alfvén wave. Whether the wave-period-averaged energy transfer to ions is positive or negative depends on the relative phases of $E_{\parallel}$ and $E_{\perp}$ to the self-consistently determined components of the ion current density $\mathbf{j}_{i}$. For the case of $\beta_{i}=3$ pictured in Fig. 7, at scales $k_{\perp}\rho_{i}\lesssim 2$ (red dotted), this relative phase $\delta\phi$ between $E_{\parallel}$ and $j_{\parallel,i}$ must fall in the range $\pi/2\leq\delta\phi\leq 3\pi/2$ (corresponding to wave growth at the expense of ion energy), while the relative phase between $E_{\perp}$ and $j_{\perp,i}$ must fall in the range $-\pi/2\leq\delta\phi\leq\pi/2$ (corresponding to wave damping and ion energy gain). Because each individual mechanism can possibly lead to negative energy transfer while the sum always leads to damping for an isotropic Maxwellian distribution, it suggests that perhaps a separation of the particle energization rates by Landau damping and transit-time damping is not particularly important in unraveling the turbulent heating of the plasma (especially since both of these mechanisms ultimately lead to energization of the parallel degree of freedom of the particles), so that only the summed effect of the Landau-resonant collisionless damping rates is important.